13.4 g of a sample of unstable hydrated salt: Na2SO4. xH20 was found to contain 6.3g of water. Determine the number of water of crystallization
1
Expert's answer
2011-06-09T10:14:32-0400
The molar mass of the salt is m(Na2SO4* xH2O) = 2*23+32+16*4 + x (2 + 16) = 142 + x*18 [g/mol]
Then let's construct the proportion:
13.4 / 6.3 = (142 + 18x)/18x
13.4 * 18x = (142 + 18x) * 6.3
241.2x = 894.6 + 113.4x
127.8x = 894.6
x = 7

Na2SO4* 7H2O